# BCTR: Bidirectional Conditioning Transformer for Scene Graph Generation

### v2

@@ 5.1 Comparisons with State-of-the-Art Methods
Results of Visual Genome The results in Tab. 1 demonstrate that BCTR outperforms other baselines in mR@ performance. By leveraging BCG and RFA, BCTR learns balanced feature representations for SGG, achieving a notable improvement of 14% and 21% in mR@K over state-of-the-art methods, indicating enhanced recall across various categories. In particular, BCTR excels in the body and tail categories, exceeding the second-best method by 18% and 67%, respectively. This performance is attributed to RFA, which constrains the feature space, enabling BCG to capture generalizable feature interactions even from rare samples, thereby boosting tail-category performance.
@@
Our R@K is lower because R@K and mR@K emphasize different performance aspects. Given the imbalanced VG dataset, R@K focuses on head predicates with abundant samples, while mR@K better reflects tail predicates, making it a fairer metric for long-tail datasets by minimizing the influence of dominant relationships [2, 5, 17]. Additionally, our DETR-based detector struggles with small entities common in the VG dataset [18], which contributes to the lower R@K. Improving small-object detection would enhance R@K. Experimental results in Tab. 1 further confirm that, when combined with statistical long-tail strategies, BCTR surpasses other methods in the body and tail categories and significantly improves mR@ performance.
Results of Open Image V6 The results of Open Image V6 are presented in Tab. 2. These experimental results demonstrate that our approach has significantly enhanced the recall performance of SGG while achieving superior or comparable weighted mAP metrics. Notably, mR@50 has improved by 6.19, demonstrating that the SGG tasks benefit from the proposed bidirectional conditioning and random feature alignment.
@@ 5.2 Ablation Studies
Table 4: Ablation Study on Feature Interactions
Method mR@20/50/100 R@20/50/100 H/B/T
BCG-UniAtt 7.6/13.1/17.5 19.6/24.2/27.3 27.0/21.3/10.8
BCG-BiAtt 7.7/13.2/17.6 19.9/24.6/27.6 27.2/21.5/10.9
Table 5: Ablation Study on Distillation Strategy
Method mR@20/50/100 R@20/50/100 H/B/T
TC 7.8/13.0/17.8 20.1/25.1/28.3 28.2/22.4/10.0
TC-RFA 7.8/13.3/18.2 20.4/25.1/28.2 28.1/22.5/10.9
LP 7.5/13.0/17.2 20.1/25.3/28.7 28.4/22.2/8.9
LP-RFA 7.5/13.0/17.6 20.2/25.2/28.3/ 28.0/21.9/10.2
@@
Model Components The results of the ablation study on model components are presented in Tab. 3, showing the performance of four model variants obtained by combining BCG and RFA. In Tab. 3, models with BCG include two additional decoder layers (i.e., bidirectional attention layers) compared to models without BCG. For fair comparison, we add two extra decoder layers to the predicate decoder in the models without BCG. The results demonstrate that both BCG and RFA contribute to improvements in mR@ and R@. Combining the two modules further improves recall in the tail category, resulting in a superior mR@100 performance of 18.4. We attribute this improvement to BCG’s ability to learn better feature interactions within a semantically aligned space, which is regularized by RFA, thereby showing superior generalization to infrequent relationship combinations (i.e., tail categories). The ablation study indicates that BCG and RFA are highly compatible, and integrating the two modules further enhances performance on the scene graph generation task.
Table 6: The Performance of Zero-shot Triplets Generation
Method zR@50/100
BGNN [17] 0.4/0.9
VCTree-TDE [29] 2.6/3.2
SGTR [18] 2.4/5.8
ISGG [12] 3.9/5.6
SG2JOIT [9] 2.5/3.7
BCTR (ours) w/o BCG+RFA 3.6/5.1
BCTR (ours) w/ BCG 3.7/5.4
BCTR (ours) w/ RFA 4.1/6.0
BCTR (ours) w/ BCG+RFA 4.4/6.2
Feature Interaction Strategy We compare the performance of two different internal message interaction patterns: Unidirectional Attention (UniAtt) and Bidirectional Attention (BiAtt). In UniAtt, only predicate-to-entity attention is activated, while in BiAtt, both entity-to-predicate and predicate-to-entity attention are enabled. The experimental results are reported in Tab. 4, where BiAtt demonstrates better performance than UniAtt, indicating the effectiveness of the designed Bidirectional Attention for the scene graph generation task.
@@
Distillation strategy We conduct ablation experiments on distillation strategies to elucidate the effectiveness of the proposed RFA. The results are presented in Tab. 5, where Trainable Classifier (TC) and Learnable Prompt (LP) [37] represent two CLIP-based classifiers. The results indicate that Random Feature Alignment enhances the performance of both classifiers. We attribute this improvement to RFA’s ability to align decoder features with CLIP, while enabling queries to capture additional objects and predicates in the image that may not be represented by CLIP. Furthermore, the results demonstrate that TC outperforms LP on scene graph generation task.
@@
Zero-shot Recall The Zero-shot Recall (zR@) of the proposed model is reported in Tab. 6. zR@ measures the recall of triplets unseen during training, reflecting the model’s generalization ability. The results show that Bidirectional Conditioning Generator provides a minor improvement in zR@, with zR@50/100 increasing by 0.1 and 0.3, respectively. When combined with Random Feature Alignment, zR@ significantly improves to 4.4 and 6.2, surpassing other baselines. This finding aligns with the observations in Tab. 3, suggesting that feature interactions learned in semantically aligned spaces generalize more effectively to unseen category relationships.
Table 7: Ablation Study on Mask Ratio
Mask Ratio mR@20/50/100 R@20/50/100 H/B/T
0 7.8/13.0/17.8 20.1/25.1/28.3 28.2/22.4/10.0
0.25 7.8/13.2/17.8 20.2/24.9/27.9 27.8/22.5/10.2
0.5 7.8/13.3/18.2 20.4/25.1/28.2 28.1/22.5/10.9
0.75 7.5/13.3/17.9 20.2/24.8/28.0/ 28.0/22.4/10.5
Figure 4: Qualitative results of our method and another method on the VG dataset. When leveraging identical DETR-based detectors, the bidirectional interaction mechanism of BCTR reduces missed detections (highlighted as yellow nodes) and enhances performance on SGG tasks.
@@
Mask Ratio We set $\alpha$ between 0 and 1 to evaluate how the mask ratio affects the model. The experimental results show high R@K values when $\alpha$ is close to 0, as the model captures the main content of the image after aligning with CLIP feature space, which accounts for a higher proportion of the VG dataset. As $\alpha$ increases, the model captures more diverse content, leading to improvements in mR@K. However, when $\alpha$ approaches 1, performance declines due to insufficient feature distillation.
Qualitative Analysis We visualize the detection results of the VG dataset in Fig. 4. For clarity, only the directed edges matching the ground-truth are shown. The experimental results indicate that, with identical DETR-based detectors, the bidirectional interaction mechanism in BCTR mutually enhances entity and predicate detection, leading to better performance on SGG compared to the baseline.
@@ 6 Conclusion
In this study, we propose a novel bidirectional conditioning factorization in semantic-aligned space for SGG and implement it by developing an end-to-end SGG model BCTR. BCTR enhances the performance of SGG tasks in two main aspects. First, the Bidirectional Conditioning Generator is designed to facilitate information interaction between entity and predicate predictions through internal bidirectional conditioning and external iterations. Second, Random Feature Alignment is introduced to facilitate BCG learning interaction patterns by randomly aligning the feature space with a pre-trained visual language model, improving generalization for semantically related relationships during inference. We perform experiments on several datasets: Visual Genome and Open Images V6, and the results demonstrate that our method achieves state-of-the-art performance.
Potential Limitation. The main limitation of the proposed method is the ongoing challenge of distilling features from VLPMs to SGG. Although RFA helps bridge the gap between CLIP’s focus on global features and SGG’s need for local predicates, further exploration of fine-grained VLPM enhancements could improve performance.

### v3

@@ 5.3.2 Results of Visual Genome
The experimental results in Tab. 1 demonstrate that the BCTR model outperforms other baselines in mR@k performance. By leveraging BCG and RFA, BCTR learns balanced feature representations for scene graph generation, achieving significant improvements of 8%, 14% and 21% in mR@k over the second-best method, indicating enhanced recall across various categories. In particular, BCTR excels in the body and tail categories, surpassing the second-best method by 18% and 68%, respectively. This performance is attributed to RFA, which constrains the feature space, enabling BCG to capture generalizable feature interactions even from rare samples, thereby boosting performance in tail categories.
@@
BCTR’s R@k appears lower because R@k and mR@k emphasize different aspects of model capability. In long-tailed datasets, the performance trade-off between head classes (categories with abundant data) and tail classes (categories with limited data) remains a central challenge [13]. Specifically, in the imbalanced VG dataset, R@k primarily reflects performance on head predicates, whereas mR@k highlights effectiveness on tail predicates. To this end, prior works [46, 31] have introduced the F@k metric to evaluate R@k and mR@k jointly. Results demonstrate that BCTR achieves superior F@k scores compared to other methods, indicating its balanced and robust performance in the SGG task. Additionally, our DETR-based detector struggles with small entities which are common in the VG dataset [20], leading to lower R@k. Improving small-object detection would enhance R@k. The experimental results in Tab. 1 further confirm that, when combined with statistical long-tail strategies, BCTR surpasses other methods in the body and tail categories and significantly improves mR@k.
5.3.3 Results of Open Image V6
The results on the Open Image V6 dataset are presented in Tab. 2. The experimental results show that BCTR achieves the highest performance across all class-balanced metrics and also gets comparable performance in R@50. In particular, BCTR outperforms the second-best method by 11% in mR@50. The F@k results indicate that BCTR achieves better comprehensive performance on the Open Image dataset than other baselines. These results indicate that random feature alignment facilitates the bidirectional conditioning generator to effectively learn generalizable feature interaction patterns from imbalanced datasets, thereby enhancing the model’s performance across various categories.
@@ 5.4 Ablation Studies
Table 3: Ablation Study on Model Components. *Parameters are computed with 6-layer entity, predicate, and indicator decoders. FPS indicates the inference speed measured on a single V100 GPU. Epoch time refers to the duration of one training epoch on 8×V100 GPUs.
BCG RFA mR@ R@ mR@100 Params*(M) FPS Epoch time (min)
20 50 100 20 50 100 Head Body Tail
7.4 12.8 17.0 19.6 24.2 27.3 26.9 21.3 9.7 96.1 5.1 25.3
7.7 13.2 17.6 19.9 24.6 27.6 27.2 21.5 10.9 99.6 3.8 33.0
7.9 13.2 17.9 20.4 25.1 28.2 28.1 22.4 10.5 96.7 4.6 31.7
8.1 13.7 18.4 20.1 24.8 27.7 27.7 22.0 11.9 100.2 3.7 38.4
Table 4: Ablation Study on Distillation Strategy.
$\text{RFA}_{\text{ENT}}$ $\text{RFA}_{\text{REL}}$ Loss mR@20 mR@50 mR100 R@20 R@50 R@100 Head Body Tail
L1 7.3 12.9 17.6 20.1 24.8 27.9 27.6 22.0 10.2
L1 7.5 12.9 17.7 20.4 25.3 28.5 28.4 22.1 10.0
L1 7.9 13.2 17.7 20.4 25.3 28.4 28.6 22.2 10.0
L2 7.8 13.3 18.1 20.4 25.2 28.3 28.3 22.5 10.2
L1 7.9 13.6 18.3 20.5 25.4 28.6 28.4 22.6 11.0
5.4.1 Model Components
@@
The results of the ablation study on model components are presented in Tab. 3, showing the performance of four model variants obtained by combining BCG and RFA. In Tab. 3, models with BCG include two additional decoder layers (i.e., bidirectional attention layers) compared to models without BCG. For a fair comparison, we add two extra decoder layers to the predicate decoder in the models without BCG. The results demonstrate that both BCG and RFA contribute to improvements in mR@k and R@k. Combining the two modules further improves recall in the tail category, resulting in a superior mR@100 performance of 18.4. We attribute this improvement to BCG’s ability to learn better feature interactions within a semantically aligned space, which is regularized by RFA, thereby improving generalization to infrequent relationship combinations (i.e., tail categories). These results indicate that BCG and RFA are highly compatible, and integrating the two modules further enhances the performance of SGG.
The impacts of RFA and BCG on the model’s computational complexity and performance are reported in Tab. 3. Specifically, BCG and RFA introduce 3.5M and 0.6M additional parameters, respectively, resulting in an overall parameter increase of approximately 4.3% compared to the baseline. During inference, BCG causes slower inference due to the added bidirectional feature interaction mechanisms, while RFA primarily functions during training and thus has negligible impact on inference speed. Regarding training performance, both components increase training time, with the overall training duration increases by 52% when BCG and RFA are employed together. In summary, the increased computational cost of BCTR stems primarily from its architectural design rather than the number of additional parameters.
5.4.2 Distillation Strategy
The analysis of the effects of different distillation strategies on model performance is presented in Tab. 4. We begin by evaluating the impact of distillation applied to different feature spaces. Specifically, we report the results of distilling only the entity features, only the relation features, and both simultaneously. The results indicate that the simultaneous distillation of both entity and relation features yields the best performance. Notably, distilling relation features alone achieves better performance than distilling entity features alone. We attribute this to the greater difficulty of relation detection compared to entity detection in the SGG task, whereby enhancing relation representations through distillation contributes more significantly to overall performance improvement.
In addition, we compare the effects of different mimic losses on distillation performance. As shown in Tab. 4, the L1 loss yields better task performance than the L2 loss. This observation is consistent with the findings reported in [24], which suggest that L1 loss is more suitable for feature distillation in relation detection tasks.
5.4.3 Classifier
@@
We conduct ablation experiments on distillation strategies to elucidate the effectiveness of the proposed RFA. The results are presented in Tab. 5, where Trainable Classifier (TC) and Learnable Prompt (LP) [49] represent two CLIP-based classifiers. TC is the classifier introduced in Section 4.3.2, which optimizes the parameters of the classifier during training. LP is the classifier introduced in [49], which optimizes learnable prompts during training and performs classification through the text encoder of CLIP. The results indicate that Random Feature Alignment enhances the performance of both classifiers, particularly in terms of tail categories and mR@k. We attribute this improvement to RFA’s ability to align decoder features with CLIP, enabling the features of tail categories to be well-represented and correctly classified, even if these categories are infrequent during training. Furthermore, the results demonstrate that TC outperforms LP in the scene graph generation task.
Table 5: Ablation Study on Classifier. Trainable Classifier (TC) and Learnable Prompt (LP) [49] represent two different types of classifiers based on CLIP features.
Method mR@20 mR@50 mR100 R@20 R@50 R@100 Head Body Tail
TC 7.8 13.0 17.8 20.1 25.1 28.3 28.2 22.4 10.0
TC-RFA 7.8 13.3 18.2 20.4 25.1 28.2 28.1 22.5 10.9
LP 7.5 13.0 17.2 20.1 25.3 28.7 28.4 22.2 8.9
LP-RFA 7.5 13.0 17.6 20.2 25.2 28.3 28.0 21.9 10.2
Table 6: Experiments on Zero-shot Triplets Generation.
Method zR@50 zR@100
BGNN [22] 0.4 0.9
VCTree-TDE [36] 2.6 3.2
SGTR [20] 2.4 5.8
ISGG [14] 3.9 5.6
BCTR (ours) w/o BCG+RFA 3.6 5.1
BCTR (ours) w/ BCG 3.7 5.4
BCTR (ours) w/ RFA 4.1 6.0
BCTR (ours) w/ BCG+RFA 4.4 6.2
5.4.4 Zero-shot Recall
@@
The Zero-shot Recall@k (zR@k) of different models is reported in Tab. 6. zR@k measures the recall of triplets that are unseen during training (i.e., combinations of subject, predicate, and object that does not appear during training), reflecting the model’s generalization performance. The results show that the Bidirectional Conditioning Generator provides a minor improvement in zR@k, with zR@50/100 increasing by 0.1 and 0.3, respectively. When combined with Random Feature Alignment, zR@k improves significantly to 4.4 and 6.2, exceeding the second-best method by 13% and 7%, respectively. This finding aligns with the observations in Tab. 3, suggesting that feature interactions learned in semantically aligned spaces generalize more effectively to unseen category relationships.
Table 7: Ablation Study on Mask Ratio.
Mask Ratio mR@20 mR@50 mR100 R@20 R@50 R@100 Head Body Tail
0 7.8 13.0 17.8 20.1 25.1 28.3 28.2 22.4 10.0
0.25 7.8 13.2 17.8 20.2 24.9 27.9 27.8 22.5 10.2
0.5 7.8 13.3 18.2 20.4 25.1 28.2 28.1 22.5 10.9
0.75 7.5 13.3 17.9 20.2 24.8 28.0 28.0 22.4 10.5
5.4.5 Mask Ratio
@@
We set the mask ratio $\alpha$ between 0 and 1 to evaluate its effect on model performance. A larger value of $\alpha$ indicates that the model has conducted more thorough feature distillation during training. In this case, the feature spaces of entities and predicates are aligned with CLIP features, but some detailed features that are not encoded by CLIP may be lost. Conversely, a smaller $\alpha$ means that the feature spaces of entities and predicates are more aligned with the dataset. The experimental results are presented in Tab. 7. The R@k is high when $\alpha$ is close to 0, as the model captures the main content of the image after aligning with the CLIP feature space, which accounts for a higher proportion of the VG dataset. As $\alpha$ increases, the model captures more diverse content, leading to improvements in mR@k. However, when $\alpha$ approaches 1, performance declines due to insufficient feature distillation. The experimental results show that the proposed model achieves the best performance when $\alpha$ is set to 0.5.
Table 8: Experiments on different image resolutions.
Reslution mR@20 mR@50 mR100 R@20 R@50 R@100 Head Body Tail
400$\times$400 7.3 12.7 16.6 18.1 22.3 24.9 25.2 20.6 9.9
500$\times$500 7.9 13.6 17.9 19.8 24.4 27.2 27.1 21.9 11.1
600$\times$600 8.1 13.7 18.4 20.1 24.8 27.7 27.7 22.0 11.9
700$\times$700 8.0 13.5 18.0 20.0 24.6 27.4 27.4 21.4 11.6
800$\times$800 8.0 13.4 17.9 19.9 24.5 27.3 27.2 21.7 11.3
Table 9: Experiments on different scenarios.
Model Scene mR@20 mR@50 R@20 R@50
SGTR [20] Simple 7.8 12.4 19.2 23.9
BCTR (ours) Simple 8.8 14.1 19.8 24.2
SGTR [20] Normal 7.4 12.8 19.9 25.0
BCTR (ours) Normal 9.0 14.1 20.5 25.4
SGTR [20] Complex 7.3 12.2 19.5 24.7
BCTR (ours) Complex 7.5 13.3 19.9 25.0
5.4.6 Image Resolution
We conducted an assessment of the trained model’s generalization across different image resolutions. Initially, the base model is trained on images with a resolution of 600$\times$600. Subsequently, the model’s performance is evaluated on images of various resolutions. The experimental results are presented in Table 8. The results indicate that performance degrades as the image resolution deviates from the baseline resolution of 600$\times$600. The performance drop is more pronounced at lower resolutions, primarily because the DETR-based detector struggles to effectively detect small objects at reduced scales. To improve model robustness to varying input image resolutions in real-world deployment scenarios, techniques such as multi-scale random scaling during training [39], multi-scale ensemble inference [32], and auxiliary super-resolution modules [47] can be employed.
5.4.7 Content Complexity
We compare the performance of BCTR across diverse scenarios. Specifically, we analyze the number of triplets within the images of the VG test dataset. Subsequently, we partitioned the test set into three equal-sized subsets based on the number of triplets in the images. The Simple scene set includes images with fewer than 4 triplets; the Normal scene set includes images with 4 to 8 triplets; and the Complex scene set includes images with more than 8 triplets. The performance of different models in various scenarios is presented in Table 9. The results demonstrate that our method consistently outperforms the baseline across varying levels of complexity, including highly complex scenarios.
5.5 Qualitative Analysis
We visualize the detection results of the VG dataset in Fig. 4. The test images cover various scenarios such as indoor, outdoor, human, animals, etc. For clarity, only the directed edges matching the ground-truth are shown. The experimental results show that BCTR detects more accurate relationships than the baseline, whether they are prominent relationships (e.g., umbrella on table, flower on table) or detailed relationships (e.g., women wearing pant, person riding bike) in the images. It is worth noting that SGTR [20] and BCTR adapt the same detector, yet BCTR achieves better results. This suggests that the bidirectional interaction mechanism of BCTR mutually enhances entity and predicate detection, leading to superior performance.
Figure 4: Qualitative results of our method and another method on the VG dataset. When leveraging identical DETR-based detectors, the bidirectional interaction mechanism of BCTR reduces missed detections (highlighted as yellow nodes) and enhances performance on the SGG task.
@@ 6 Conclusion
In this study, we propose a novel bidirectional conditioning factorization in semantic-aligned space for SGG and implement it by developing an end-to-end SGG model BCTR. BCTR enhances the performance of SGG tasks in two main aspects. First, the Bidirectional Conditioning Generator is designed to facilitate information interaction between entity and predicate predictions through internal bidirectional conditioning and external iterations. Second, Random Feature Alignment is introduced to facilitate BCG in learning interaction patterns by randomly aligning the feature space with a pre-trained visual language model, improving generalization for semantically related relationships during inference. We perform experiments on several datasets: Visual Genome and Open Images V6, and the results demonstrate that our method achieves state-of-the-art performance.
The limitation of the proposed method lies in the ongoing challenges associated with distilling features from VLPMs to SGG. Firstly, while RFA is introduced to address the discrepancy between CLIP’s focus on capturing global features and the SGG task’s focus on local predicates, exploring more fine-grained VLPM enhancements could be beneficial for improving performance in the future. Additionally, CLIP exhibits poor performance in capturing structured knowledge, such as object properties and predicates, which are crucial components in SGG tasks. Therefore, exploring feature distillation with more advanced CLIP models is an interesting direction for further investigation.
Acknowledgments
This work was supported in part by the National Key Research and Development Program of China under Grant 2023YFB4705000, in part by the National Natural Science Foundation of China under 62303455, 62373039, U23B2038, U24A20282, 62273342, 62403461, and 62122087, in part by Beijing Natural Science Foundation under Grant L233006, in part by the Youth Program and the Open Projects Program of State Key Laboratory of Multimodal Artificial Intelligence Systems.